3) Find the Laplace transform of 1-e^{-2t}/t


1
Expert's answer
2021-09-17T02:36:07-0400


here F(ω)F(ω) is Laplace transform of f(t)f(t). We choose f(t)=(1−e−t)f(t)=(1−e−t) and it is easy to show that

F(s)=L[1−e−t]=1s−1s+1F(s)=L[1−e−t]=1s−1s+1